# LiveCodeBench: Holistic and Contamination Free Evaluation of Large Language Models for Code

### v1

@@ 5.2 Performance and Model Comparisons
Holistic Evaluations. We have evaluated the models across the four scenarios currently available in LiveCodeBench. Figure 2 displays the performance of models on all scenarios along the axes of the polar chart. First, we observe that the relative order of models remains mostly consistent across the scenarios. This is also supported by high correlations between Pass@1 metric across the scenarios – over across all pairs as shown in Figure 13. Interestingly, the correlations are larger for related tasks, for generation and self-repair, and for test output prediction and code execution. This correlation drops to for generation and test prediction scenarios.
@@
Indeed, HumanEval is an easier benchmark with small and isolated programming problems and thus easier to overfit on. In contrast, LiveCodeBench problems are sourced from reputable coding platforms offering more challenging problems with higher diversity and difficulty levels. This potential overfitting is particularly exemplified by DS-Ins-1.3B which achieves % Pass@1 on HumanEval+ but only % on LCB-Easy. Thus, while it boasts better performance compared with Gemini-Pro and Claude-Ins-1 on HumanEval+, it performs considerably worse on LCB-Easy. Similarly, DS-Ins-6.7B and MC-6.7B perform better than Mistral-L, Claude-2, and Claude-3-Sonnet on HumanEval+ but are considerably worse on LCB-Easy.
Highlighting the gap between SoTA and open models. One distinct observation from our evaluations is the large gap between SoTA models and open models across all scenarios. Particularly, GPT-4-Turbo, GPT-4, and Claude-3-Opus lead across the benchmarks with wide performance margins over other models. This distinguished LiveCodeBench from prior benchmarks (like HumanEval) where various open models have achieved similar or better performance. For example, DS-Ins-33B is merely point behind GPT-4-Turbo on HumanEval+ but points (%) on LCB code generation scenario. This gap either holds or sometimes even amplifies across other scenarios. For instance, consider test output prediction and code execution (with COT) where GPT-4-Turbo leads the DS-Ins-33B model by and respectively!
@@
Comparing Base Models. We use three families of base models – DeepSeek, CodeLLaMa, and StarCoder2 and compare them on the code generation scenario. A one-shot prompt is used for all models to avoid any formatting and answer extraction issues. We find DS models are significantly better than both CodeLLaMa and StarCoder2 base models with a DS-Base-6.7B model even outperforming both CL-Base-34B and SC2-Base-15B models. Next, we observe that SC2-Base-15B also outperforms the CL-Base-34B model (similiar to findings in Lozhkov et al. (2024)). Note that these differences can potentially be attributed to data curation approaches. For instance, StarCoder2 models (and potentially DeepSeeks as discussed in Section 5.1) use competition problems in the pre-training corpus highlighting the need for diverse and clean data for training code LLMs. Similarly, training methodologies (like epochs and learning schedules) can also play a role in the difference and we leave a thorough investigation of this for future work.
Role of Instruction Tuning. We find that fine-tuning improves performance on both HumanEval+ and LiveCodeBench for the code generation scenario. Particularly, DS-Ins-33B and Phind-34B achieve and Pass@1 on LiveCodeBench improving over their base models by and points respectively. Similar gains are observed in previous benchmarks (like HumanEval+) as well. This highlights the importance of good instruction datasets for building strong code LLMs. Thus, even though fine-tuning is assumed to not significantly alter model capabilities, it is essential for improving performance by eliciting the right model behavior.
At the same time, we note that the base models have aligned performances on LCB code generation and HumanEval+ benchmarks and lie within or close to the green shaded region in Figure 5. However, the fine-tuned models exhibit a large performance gap, with much better performance on HumanEval+. On the other hand, the closed-access models are still aligned across both benchmarks. This suggests that the fine-tuning data for open models might not be as diverse as that for closed models, leading to a lack of generalization to different kinds of problems.
Comparing open-access instruction-tuned models. Here, we compare various fine-tuned variants of the DeepSeek and CodeLLaMa base models across different model sizes. We find that fine-tuned DeepSeek models lead in performance, followed by Phind-34B and CL-Base-{7, 13, 34}B models across most scenarios. Broadly, we find that model performances correlate with model sizes. For example,Phind-34B model outperforms the B models across all scenarios.
@@
Open-Access vs Closed-Access Models. In general, closed (API) access model families generally outperform the open access models. The gap is only closed by three models, namely DS-Ins-33B, WC-33B, and Phind-34B which reach the performance levels of the closed models. For instance, in the code generation scenario (Figure 2 right), these models reach close to or even outperform closed access models like Gemini-Pro, GPT-3.5-turbo, and Claude-3-Sonnet. The performances vary across scenarios with the closed-access models performing better in test output prediction and code execution scenarios. Overall, our findings confirm that a combination of strong base models and high-quality fine-tuning datasets is a viable recipe for good code LLMs.
@@ 6.1 Code Generation
Language Models for Code Generation. Starting with Codex (Chen et al., 2021), there are over a dozen code LLMs. These include CodeT5 (Wang et al., 2021, 2023), CodeGen (Nijkamp et al., 2022), SantaCoder (Allal et al., 2023), StarCoder (Li et al., 2023b), AlphaCode (Li et al., 2022), InCoder (Fried et al., 2022), and CodeGeeX (Zheng et al., 2023). As of February 2024, DeepSeek (Bi et al., 2024), StarCoder Lozhkov et al. (2024); Li et al. (2023b) and CodeLLaMa (Roziere et al., 2023) are the most popular open models. Many downstream models resulted from fine-tuning them on synthetically generated data, such as WizardCoder (Luo et al., 2023), MagiCoders (Wei et al., 2023b), and Phind-34B.
@@
A few benchmarks specifically measure competitive programming, such as APPS (Hendrycks et al., 2021), CodeContests (Li et al., 2022), xCodeEval (Khan et al., 2023), and LeetCode-Hard (Shinn et al., 2023), and TACO (Li et al., 2023c). Methods such as AlphaCode (Li et al., 2022), AlphaCode 2(Gemini Team et al., 2023), ALGO (Zhang et al., 2023d), Parsel (Zelikman et al., 2022), code cleaning (Jain et al., 2023), code explainations (Li et al., 2023a), analogical reasoning (Yasunaga et al., 2023), and AlphaCodium (Ridnik et al., 2024) have been pushing the boundaries of what is possible with LLMs in this domain. The biggest differentiating factor between LiveCodeBench and these benchmarks is that our benchmark is continuously updated and contains more tasks such as code repair, code execution, and test output prediction capturing more facets of code reasoning and code generation.
@@ 6.3 Contamination
Data contamination and test-case leakage have received considerable attention Oren et al. (2024); Golchin and Surdeanu (2023); Weller et al. (2023); Roberts et al. (2024) as LLMs might be getting trained on benchmarks. Sainz et al. (2023) demonstrated contamination by simply prompting the model to highlight its contamination. Some detection methods have also been built to avoid these cases (Shi et al., 2023; Zhou et al., 2023). For code, Riddell et al. (2024) use edit distance and AST-based semantic-similarity to detect contamination. Our work addresses contamination by live updates.
@@ C.1 Models
We describe the details of models considered in our study in Table 2.
@@
deepseek-ai/deepseek-coder-33b-instruct |
DS-33b-Ins |
2023-07-30 |
deepseek-ai/deepseek-coder-6.7b-instruct |
DS-6.7b-Ins |
2023-07-30 |
deepseek-ai/deepseek-coder-1.3b-instruct |
DS-1.3b-Ins |
2023-07-30 |
codellama/CodeLlama-34b-Instruct-hf |
CL-34b-Ins |
2023-01-01 |
codellama/CodeLlama-13b-Instruct-hf |
CL-13b-Ins |
2023-01-01 |
codellama/CodeLlama-7b-Instruct-hf |
CL-7b-Ins |
2023-01-01 |
WizardLM/WizardCoder-Python-34B |
WC-34B |
2023-01-01 |
WizardLM/WizardCoder-Python-33B-V1.1 |
WC-33B |
2023-07-30 |
Phind/Phind-CodeLlama-34B-v2 |
Phind-34B |
2023-01-01 |
ise-uiuc/Magicoder-S-DS-6.7B |
MC-6.7B |
2023-07-30 |
ise-uiuc/Magicoder-S-CL-7B |
MC-7B |
2023-01-01 |
gpt-3.5-turbo-0301 |
GPT-3.5-Turbo-0301 |
2021-10-01 |
gpt-3.5-turbo-1106 |
GPT-3.5-Turbo-1106 |
2021-10-01 |
gpt-3.5-turbo-0613 |
GPT-3.5-Turbo-0613 |
2021-10-01 |
gpt-4-0314 |
GPT-4-0314 |
2021-10-01 |
gpt-4-0613 |
GPT-4-0613 |
2021-10-01 |
gpt-4-1106-preview |
GPT-4-Turbo-1106 |
2023-04-30 |
claude-2 |
Claude-2 |
2022-12-31 |
claude-instant-1 |
Claude-Instant-1 |
2022-12-31 |
claude-3-opus |
Claude3-O |
2023-07-30 |
claude-3-sonnet |
Claude3-S |
2023-07-30 |
gemini-pro |
Google-Gemini-Pro |
mistral-large-latest |
Mistral-L |
C.2 Code Generation
Below we provide the prompt format (with appropriate variants adding special tokens accommodating each instruct-tuned model) used for this scenario.
C.3 Self Repair
Below we provide the prompt format (with appropriate variants adding special tokens accommodating each instruct-tuned model) used for this scenario.
C.4 Code Execution
Below we provide the prompts for code execution with and without CoT. The prompts are modified versions of those from (Gu et al., 2024) to fit the format of the samples in our benchmark.
C.5 Test Output Prediction
Below we provide the prompt format (with appropriate variants adding special tokens accommodating each instruct-tuned model) used for this scenario.
Appendix D Results
D.1 Contamination
Figure 10 demonstrates contamination in DeepSeek in self repair and test output prediction scenarios.
D.2 All Results
Below we provide the tables comprising of results across different LiveCodeBench scenarios.
| Model Name | Easy | Medium | Hard | Total |
| Claude-2 | 61.5 | 8.0 | 0.4 | 23.3 |
| Claude-3-Opus | 76.5 | 14.7 | 6.4 | 32.5 |
| Claude-3-Sonnet | 63.3 | 8.3 | 1.1 | 24.2 |
| Claude-Instant-1 | 60.1 | 6.4 | 1.8 | 22.8 |
| Cllama-13b-Ins | 37.4 | 2.7 | 0.0 | 13.4 |
| Cllama-34b-Ins | 30.4 | 3.4 | 1.8 | 11.8 |
| Cllama-7b-Ins | 31.3 | 3.1 | 0.0 | 11.5 |
| CodeLlama-13b-Base | 25.2 | 1.3 | 0.0 | 8.8 |
| CodeLlama-34b-Base | 32.7 | 2.7 | 0.2 | 11.8 |
| CodeLlama-7b-Base | 20.4 | 0.7 | 0.0 | 7.0 |
| DSCoder-1.3b-Base | 16.4 | 0.9 | 0.0 | 5.8 |
| DSCoder-1.3b-Ins | 26.4 | 2.1 | 0.0 | 9.5 |
| DSCoder-33b-Base | 39.4 | 2.3 | 0.0 | 13.9 |
| DSCoder-33b-Ins | 56.5 | 7.1 | 1.1 | 21.6 |
| DSCoder-6.7b-Base | 36.9 | 1.2 | 0.2 | 12.8 |
| DSCoder-6.7b-Ins | 46.8 | 5.3 | 1.1 | 17.7 |
| GPT-3.5-Turbo-0125 | 52.9 | 9.8 | 0.0 | 20.9 |
| GPT-3.5-Turbo-0301 | 52.1 | 9.6 | 0.4 | 20.7 |
| GPT-4-0613 | 74.6 | 21.3 | 0.5 | 32.2 |
| GPT-4-Turbo-1106 | 81.9 | 24.5 | 1.1 | 35.8 |
| Google-Gemini-Pro | 50.5 | 8.5 | 0.0 | 19.6 |
| MagiCoderS-CL-7B | 33.8 | 2.8 | 0.0 | 12.2 |
| MagiCoderS-DS-6.7B | 48.0 | 7.9 | 0.0 | 18.6 |
| Mistral-Large | 58.5 | 13.0 | 1.1 | 24.2 |
| Phind-34B-V2 | 55.4 | 6.0 | 0.2 | 20.5 |
| StarCoder2-15b | 37.1 | 3.2 | 0.0 | 13.4 |
| StarCoder2-3b | 27.5 | 0.9 | 0.0 | 9.5 |
| StarCoder2-7b | 33.5 | 1.7 | 0.0 | 11.8 |
| WCoder-33B-V1.1 | 56.4 | 8.1 | 0.4 | 21.6 |
| WCoder-34B-V1 | 45.8 | 4.2 | 0.4 | 16.8 |
| Model Name | Easy | Medium | Hard | Total |
| Claude-2 | 66.2 | 10.3 | 0.4 | 25.6 |
| Claude-3-Opus | 83.1 | 23.7 | 6.7 | 37.8 |
| Claude-3-Sonnet | 72.6 | 11.8 | 2.2 | 28.9 |
| Claude-Instant-1 | 64.4 | 7.1 | 2.2 | 24.6 |
| Cllama-13b-Ins | 43.1 | 3.0 | 0.0 | 15.3 |
| Cllama-34b-Ins | 31.5 | 3.5 | 1.8 | 12.3 |
| Cllama-7b-Ins | 31.9 | 3.1 | 1.5 | 12.1 |
| DSCoder-1.3b-Ins | 29.5 | 2.1 | 0.0 | 10.6 |
| DSCoder-33b-Ins | 60.7 | 8.1 | 1.5 | 23.4 |
| DSCoder-6.7b-Ins | 49.9 | 5.7 | 1.1 | 18.9 |
| GPT-3.5-Turbo-0125 | 59.3 | 11.9 | 0.5 | 23.9 |
| GPT-3.5-Turbo-0301 | 58.4 | 11.6 | 0.7 | 23.6 |
| GPT-4-0613 | 79.3 | 25.0 | 2.4 | 35.6 |
| GPT-4-Turbo-1106 | 86.9 | 36.9 | 4.0 | 42.6 |
| Google-Gemini-Pro | 53.8 | 9.4 | 0.2 | 21.1 |
| MagiCoderS-CL-7B | 36.5 | 3.1 | 0.0 | 13.2 |
| MagiCoderS-DS-6.7B | 50.6 | 8.6 | 0.0 | 19.7 |
| Mistral-Large | 71.2 | 15.6 | 3.6 | 30.1 |
| Phind-34B-V2 | 62.0 | 6.5 | 0.9 | 23.1 |
| WCoder-33B-V1.1 | 61.8 | 10.0 | 1.8 | 24.5 |
| WCoder-34B-V1 | 50.0 | 5.1 | 1.1 | 18.7 |
| Model Name | Pass@1 |
| Claude-2 | 32.7 |
| Claude-3-Opus | 58.7 |
| Claude-3-Sonnet | 34.1 |
| Claude-Instant-1 | 25.4 |
| Cllama-13b-Ins | 24.4 |
| Cllama-34b-Ins | 23.0 |
| Cllama-7b-Ins | 15.3 |
| DSCoder-1.3b-Ins | 12.5 |
| DSCoder-33b-Ins | 28.3 |
| DSCoder-6.7b-Ins | 26.5 |
| GPT-3.5-Turbo-0125 | 35.4 |
| GPT-3.5-Turbo-0301 | 32.5 |
| GPT-4-0613 | 52.9 |
| GPT-4-Turbo-1106 | 55.7 |
| Google-Gemini-Pro | 29.5 |
| MagiCoderS-CL-7B | 21.3 |
| MagiCoderS-DS-6.7B | 27.1 |
| Mistral-Large | 46.5 |
| Phind-34B-V2 | 27.2 |
| WCoder-33B-V1.1 | 26.7 |
| WCoder-34B-V1 | 24.4 |
| Model Name | Pass@1 | Pass@1 (COT) |
| Claude-2 | 31.0 | 45.0 |
| Claude-3-Opus | 36.3 | 80.9 |
| Claude-3-Sonnet | 30.5 | 39.8 |
| Claude-Instant-1 | 20.5 | 35.6 |
| Cllama-13b-Ins | 23.9 | 13.9 |
| Cllama-34b-Ins | 28.8 | 24.9 |
| Cllama-7b-Ins | 20.1 | 14.1 |
| DSCoder-1.3b-Base | 19.1 | 13.5 |
| DSCoder-1.3b-Ins | 18.8 | 17.6 |
| DSCoder-33b-Base | 29.8 | 29.0 |
| DSCoder-33b-Ins | 26.5 | 31.9 |
| DSCoder-6.7b-Base | 23.4 | 25.0 |
| DSCoder-6.7b-Ins | 23.2 | 24.0 |
| GPT-3.5-Turbo-0301 | 34.4 | 36.2 |
| GPT-4-0613 | 44.9 | 65.7 |
| GPT-4-Turbo-1106 | 40.2 | 84.7 |
| Google-Gemini-Pro | 26.6 | 37.5 |
| Mistral-Large | 37.0 | 54.8 |
Appendix E Qualitative Examples
E.1 Code Execution
We show 5 examples from the code execution task that GPT-4 (gpt-4-1106-preview) still struggles to execute, even with CoT.

### v2

@@ 5.2 Performance and Model Comparisons
Holistic Evaluations. We have evaluated the models across the four scenarios currently available in LiveCodeBench. Figure 2 displays the performance of models on all scenarios along the axes of the polar chart. First, we observe that the relative order of models remains mostly consistent across the scenarios. This is also supported by high correlations between Pass@1 metric across the scenarios – over across all pairs as shown in Figure LABEL:app:fig:correlation_matrix. Interestingly, the correlations are larger for related tasks, for generation and self-repair, and for test output prediction and code execution. This correlation drops to for generation and execution scenarios.
@@
Indeed, HumanEval is an easier benchmark with small and isolated programming problems and thus easier to overfit on. In contrast, LiveCodeBench problems are sourced from reputable coding platforms offering more challenging problems with higher diversity and difficulty levels. This potential overfitting is particularly exemplified by DS-Ins-1.3B which achieves % Pass@1 on HumanEval+ but only % on LCB-Easy. Thus, while it boasts better performance compared with Gemini-Pro and Claude-Ins-1 on HumanEval+, it performs considerably worse on LCB-Easy. Similarly, CodeQwen, DS-Ins-6.7B, and MC-6.7B perform better than Mistral-L, Claude-2, and Claude-3-Sonnet on HumanEval+ but are considerably worse on LCB-Easy.
Highlighting the gap between SoTA and open models. One distinct observation from our evaluations is the large gap between SoTA models and open models across all scenarios. Particularly, GPT-4-Turbo, GPT-4, Gemini-Pro-1.5 and Claude-3-Opus lead across the benchmarks with wide performance margins over other models. This distinguishes LiveCodeBench from prior benchmarks (like HumanEval) where various open models have achieved similar or better performance. For example, DS-Ins-33B is merely point behind GPT-4-Turbo on HumanEval+ but points (%) on LCB code generation scenario. This gap either holds or sometimes even amplifies across other scenarios. For instance, consider test output prediction and code execution (with COT) where GPT-4-Turbo leads the DS-Ins-33B model by and respectively!
@@
Comparing Base Models. We use four families of base models – L3-Base, DeepSeek, CodeLLaMa, and StarCoder2 and compare them on the code generation scenario. A one-shot prompt is used for all models to avoid any formatting and answer extraction issues. We find L3-Base and DS models are significantly better than both CodeLLaMa and StarCoder2 base models with a DS-Base-6.7B model even outperforming both CL-Base-34B and SC2-Base-15B models. Next, we observe that SC2-Base-15B also outperforms the CL-Base-34B model (similiar to findings in Lozhkov et al. (2024)). Note that some LiveCodeBench specific differences can potentially be attributed to data curation approaches. For instance, StarCoder2 models (and potentially DeepSeeks as discussed in Section 5.1) use competition problems in the pre-training corpus.
Role of Post Training. We find that post-training improves performance on both HumanEval+ and LiveCodeBench for the code generation scenario. Particularly, L3-Ins-70B, DS-Ins-33B and Phind-34B achieve , , and Pass@1 on LCB improving over their base models by , and points respectively. Similar gains are observed in previous benchmarks (like HumanEval+) as well. This highlights the importance of good post-training datasets for building strong LLMs.
At the same time, we note that the base models have aligned performances on LCB code generation and HumanEval+ benchmarks and lie within or close to the green shaded region in Figure 5. However, the fine-tuned open models exhibit a larger performance gap, with much better performances on HumanEval+. On the other hand, the closed-access models are still aligned across both benchmarks. This suggests that the fine-tuning data for open models might not be as diverse as that for closed models, leading to a lack of generalization to different kinds of problems.
Comparing open-access instruction-tuned models. Here, we compare various fine-tuned variants of the L3-Base, DeepSeek and CodeLLaMa base models across different model sizes. We find that fine-tuned L3-Base and DeepSeek models lead in performance, followed by Phind-34B and CodeLLaMa models across most scenarios. Broadly, we find that model performances correlate with model sizes. For example, Phind-34B model outperforms the B models across all scenarios.
@@
Open-Access vs Closed-Access Models. In general, closed (API) access model families generally outperform the open access models. The gap is only closed by three models, namely L3-Ins-70B, Mixtral, and DS-Ins-33B which reach the performance levels of the closed models. For instance, in the code generation scenario (Figure 2 right), these models reach close to or even outperform closed access models like Gemini-Pro, GPT-3.5-turbo, and Claude-3-Sonnet. The performances vary across scenarios with the closed-access models performing better in test output prediction and code execution scenarios. Overall, our findings confirm that a combination of strong base models and high-quality post-training datasets is a viable recipe for good code LLMs.
@@ 6.1 Code Generation
Language Models for Code Generation. Starting with Codex (Chen et al., 2021), there are over a dozen code LLMs. These include CodeT5 (Wang et al., 2021, 2023), CodeGen (Nijkamp et al., 2022), SantaCoder (Allal et al., 2023), StarCoder (Li et al., 2023b), AlphaCode (Li et al., 2022), InCoder (Fried et al., 2022), and CodeGeeX (Zheng et al., 2023). As of May 2024, L3-Base and DeepSeek (Bi et al., 2024), StarCoder Lozhkov et al. (2024); Li et al. (2023b) and CodeLLaMa (Roziere et al., 2023) are the most popular open models. Many downstream models resulted from fine-tuning them on synthetically generated data, such as WizardCoder (Luo et al., 2023), MagiCoders (Wei et al., 2023b), and Phind-34B.
@@
A few benchmarks specifically measure competitive programming, such as APPS (Hendrycks et al., 2021), CodeContests (Li et al., 2022), CodeScope (Yan et al., 2023), xCodeEval (Khan et al., 2023), and LeetCode-Hard (Shinn et al., 2023), and TACO (Li et al., 2023c). Methods such as AlphaCode (Li et al., 2022), AlphaCode 2(Gemini Team et al., 2023), ALGO (Zhang et al., 2023d), Parsel (Zelikman et al., 2022), code cleaning (Jain et al., 2023), code explainations (Li et al., 2023a), analogical reasoning (Yasunaga et al., 2023), and AlphaCodium (Ridnik et al., 2024) have been pushing the boundaries of what is possible with LLMs in this domain. The biggest differentiating factor between LiveCodeBench and these benchmarks is that our benchmark is continuously updated, problem curation with balanced difficulty, higher tests and problem quality, and contains more scenarios such as code repair, code execution, and test output prediction capturing more facets for building agentic coding systems.
@@ 6.3 Contamination
Data contamination and test-case leakage have received considerable attention Oren et al. (2024); Golchin and Surdeanu (2023); Weller et al. (2023); Roberts et al. (2024) as LLMs might be getting trained on benchmarks. Sainz et al. (2023) demonstrated contamination by simply prompting the model to highlight its contamination. Some detection methods have also been built to avoid these cases (Shi et al., 2023; Zhou et al., 2023). For code, Riddell et al. (2024) use edit distance and AST-based semantic-similarity to detect contamination.
@@ C.1 Models
We describe the details of models considered in our study in Table LABEL:tab:models_list.
@@
| deepseek-ai/deepseek-coder-33b-instruct | DSCoder-33b-Ins | 08/30/2023 | deepseek-coder-33b-instruct |
| deepseek-ai/deepseek-coder-6.7b-instruct | DSCoder-6.7b-Ins | 08/30/2023 | deepseek-coder-6.7b-instruct |
| deepseek-ai/deepseek-coder-1.3b-instruct | DSCoder-1.3b-Ins | 08/30/2023 | deepseek-coder-1.3b-instruct |
| codellama/CodeLlama-70b-Instruct-hf | CodeLlama-70b-Ins | 01/01/2023 | CodeLlama-70b-Instruct-hf |
| openbmb/Eurus-70b-sft | Eurus-70B-SFT (n=1) | 01/01/2023 | Eurus-70b-sft |
| openbmb/Eurux-8x22b-nca | Eurux-8x22b-NCA (n=1) | 04/30/2023 | Eurux-8x22b-nca |
| codellama/CodeLlama-34b-Instruct-hf | CodeLlama-34b-Ins | 01/01/2023 | CodeLlama-34b-Instruct-hf |
| codellama/CodeLlama-13b-Instruct-hf | CodeLlama-13b-Ins | 01/01/2023 | CodeLlama-13b-Instruct-hf |
| codellama/CodeLlama-7b-Instruct-hf | CodeLlama-7b-Ins | 01/01/2023 | CodeLlama-7b-Instruct-hf |
| meta-llama/Meta-Llama-3-8B-Instruct | LLama3-8b-Ins | 01/01/2023 | Meta-Llama-3-8B-Instruct |
| meta-llama/Meta-Llama-3-70B-Instruct | LLama3-70b-Ins | 01/01/2023 | Meta-Llama-3-70B-Instruct |
| Phind/Phind-CodeLlama-34B-v2 | Phind-34B-V2 | 01/01/2023 | Phind-CodeLlama-34B-v2 |
| Smaug-2-72B | Smaug-2-72B | 01/01/2023 | Smaug-2-72B |
| Qwen-1.5-72B-Chat | Qwen-1.5-72B-Chat | 01/01/2023 | Qwen-1.5-72B-Chat |
| Qwen/CodeQwen1.5-7B | CodeQwen15-7B | 08/30/2023 | CodeQwen1.5-7B |
| Qwen/CodeQwen1.5-7B-Chat | CodeQwen15-7B-Chat | 08/30/2023 | CodeQwen1.5-7B-Chat |
| gpt-3.5-turbo-0301 | GPT-3.5-Turbo-0301 | 10/01/2021 | gpt-3.5-turbo-0301 |
| gpt-3.5-turbo-0125 | GPT-3.5-Turbo-0125 | 10/01/2021 | gpt-3.5-turbo-0125 |
| gpt-4-0613 | GPT-4-0613 | 10/01/2021 | gpt-4-0613 |
| gpt-4-1106-preview | GPT-4-Turbo-1106 | 04/30/2023 | gpt-4-1106-preview |
| gpt-4-turbo-2024-04-09 | GPT-4-Turbo-2024-04-09 | 04/30/2023 | gpt-4-turbo-2024-04-09 |
| gpt-4o-2024-05-13 | GPT-4O-2024-05-13 | 10/30/2023 | gpt-4o-2024-05-13 |
| claude-2 | Claude-2 | 12/31/2022 | claude-2 |
| claude-instant-1 | Claude-Instant-1 | 12/31/2022 | claude-instant-1 |
| claude-3-opus-20240229 | Claude-3-Opus | 04/30/2023 | claude-3-opus-20240229 |
| claude-3-sonnet-20240229 | Claude-3-Sonnet | 04/30/2023 | claude-3-sonnet-20240229 |
| claude-3-haiku-20240307 | Claude-3-Haiku | 04/30/2023 | claude-3-haiku-20240307 |
| codestral-latest | Codestral-Latest | 01/31/2024 | codestral-latest |
| gemini-pro | Gemini-Pro | 04/30/2023 | gemini-pro |
| gemini-1.5-pro-latest | Gemini-Pro-1.5-May | 04/30/2023 | gemini-1.5-pro-latest |
| gemini-1.5-flash-latest | Gemini-Flash-1.5-May | 04/30/2023 | gemini-1.5-flash-latest |
| ise-uiuc/Magicoder-S-DS-6.7B | MagiCoderS-DS-6.7B | 08/30/2023 | Magicoder-S-DS-6.7B |
| ise-uiuc/Magicoder-S-CL-7B | MagiCoderS-CL-7B | 01/01/2023 | Magicoder-S-CL-7B |
| bigcode/starcoder2-3b | StarCoder2-3b | 01/01/2023 | starcoder2-3b |
| bigcode/starcoder2-7b | StarCoder2-7b | 01/01/2023 | starcoder2-7b |
| bigcode/starcoder2-15b | StarCoder2-15b | 01/01/2023 | starcoder2-15b |
| codellama/CodeLlama-70b-hf | CodeLlama-70b-Base | 01/01/2023 | CodeLlama-70b-hf |
| codellama/CodeLlama-34b-hf | CodeLlama-34b-Base | 01/01/2023 | CodeLlama-34b-hf |
| codellama/CodeLlama-13b-hf | CodeLlama-13b-Base | 01/01/2023 | CodeLlama-13b-hf |
| codellama/CodeLlama-7b-hf | CodeLlama-7b-Base | 01/01/2023 | CodeLlama-7b-hf |
| deepseek-ai/deepseek-coder-33b-base | DSCoder-33b-Base | 08/30/2023 | deepseek-coder-33b-base |
| deepseek-ai/deepseek-coder-6.7b-base | DSCoder-6.7b-Base | 08/30/2023 | deepseek-coder-6.7b-base |
| deepseek-ai/deepseek-coder-1.3b-base | DSCoder-1.3b-Base | 08/30/2023 | deepseek-coder-1.3b-base |
| google/codegemma-7b | CodeGemma-7b-Base | 01/01/2023 | codegemma-7b |
| google/codegemma-2b | CodeGemma-2b-Base | 01/01/2023 | codegemma-2b |
| google/gemma-7b | Gemma-7b-Base | 01/01/2023 | gemma-7b |
| google/gemma-2b | Gemma-2b-Base | 01/01/2023 | gemma-2b |
| meta-llama/Meta-Llama-3-70B | LLama3-70b-Base | 01/01/2023 | Meta-Llama-3-70B |
| meta-llama/Meta-Llama-3-8B | LLama3-8b-Base | 01/01/2023 | Meta-Llama-3-8B |
| mistral-large-latest | Mistral-Large | 01/01/2023 | mistral-large-latest |
| open-mixtral-8x22b | Mixtral-8x22B-Ins | 01/01/2023 | open-mixtral-8x22b |
| open-mixtral-8x7b | Mixtral-8x7B-Ins | 01/01/2023 | open-mixtral-8x7b |
| m-a-p/OpenCodeInterpreter-DS-33B | OC-DS-33B | 08/30/2023 | OpenCodeInterpreter-DS-33B |
| m-a-p/OpenCodeInterpreter-DS-6.7B | OC-DS-6.7B | 08/30/2023 | OpenCodeInterpreter-DS-6.7B |
| m-a-p/OpenCodeInterpreter-DS-1.3B | OC-DS-1.3B | 08/30/2023 | OpenCodeInterpreter-DS-1.3B |
| command-r | Command-R | 01/01/2023 | command-r |
| command-r+ | Command-R+ | 01/01/2023 | command-r+ |